### v1

# A reduced-order framework for temperature estimation in food freezing from optimally located sensors, including turbulent conjugate flow scenarios

Abstract
This article proposes a framework for estimating temperature fields in food-freezing applications that significantly reduces computational load while ensuring accurate temperature monitoring, representing a promising technological tool for optimizing and controlling food engineering processes. The strategy is based on (i) a mathematical model of a convection-dominated problem coupling thermal convection and turbulence and (ii) a least-squares approach for solving the inverse data assimilation problem, regularized by projecting the governing dynamics onto a reduced-order model (ROM). The unsteady freezing process considers an idealized salmon slice in a freezer cabinet, modeled with temperature-dependent thermophysical properties. The forward problem is approximated using a third-order WENO finite volume solver, including an optimized second-order backward scheme for time discretization. We employ our data assimilation framework to reconstruct the temperature field from a limited number of sensor data and to estimate temperature distributions within frozen food. Sensor placement is optimized using a new greedy algorithm, relying on maximizing the observability of the reduced-order dynamics for a fixed set of sensors. The proposed approach allows efficient extrapolation from external sensor measurements to the internal temperature of the food, which is crucial for maintaining food quality.
keywords:
Food Freezing , Data assimilation , Greedy Algorithm , Reduced Order Modelling , Phase Change , Optimal Sensor Placement
1 Introduction
Freezing processes, crucial for transporting and preserving food [1, 2], have important economic implications and rely on large amounts of energy required to satisfy safe freezing conditions. Food freezing involves a highly nonlinear liquid-solid phase-change that produces significant changes in the processed food and strongly depends on the operational conditions imposed, for example, in a freezing tunnel. Liquid-solid phase change implies a boundary evolution scenario in which the phases interact, natural convection occurs, the thermophysical properties of the food change strongly, and turbulence develops in the airflow that freezes the food [3, 4]. In addition, food humidity, texture, color, and flavor depend on the evolution of the freezing rate [5, 6, 7]. All these factors influence food quality and the possibility of estimating in advance how the food temperature changes during freezing is of both scientific and technological interest.
Computational fluid dynamics has become a standard tool to support the understanding of the physics of fluids and of the heat-mass transfer that occurs in food processing. For example, computer simulations have been successfully used for evaluating convective coefficients [8, 9, 10], for estimating freezing times [11, 12, 13], for assessing food positioning for better freezing [3], and for testing the effectivity of different operational conditions [14], thereby improving the overall thermal processes. Numerical methods used for these simulations have been based on finite difference [15, 16], finite element [17, 18, 19] or finite volume methods [20, 21, 22]. The advantages and disadvantages of these numerical methodologies have been recently discussed in [23, 24].
Inverse problems in the food industry have gained attention in recent years, primarily focusing on parameter estimation techniques. Recent studies, such as those in [25], emphasize the use of inverse strategies, particularly in identifying material properties and diffusivity parameters in food products. For example, research in [26] highlights the estimation of moisture diffusivity by inverting a finite element model applied to various solid foods such as salami, biscuits, and flatbread by minimizing the distance between the numerical model and experimental results through the Levenberg–Marquardt algorithm. Just as ensuring the properties of heat-treated foods is of scientific and technological interest, efficient energy consumption in the processes is also an aspect that may be optimized using similar techniques [27]. In [28], the authors highlight that using the bootstrap technique can estimate both specific heat and thermal conductivity in a freezing range, comparing the obtained results with experimental data.
In this work, we address the estimation of the temperature fields in a freezing salmon slice from measurement data, focusing on the case of partial and low-resolution observations. This type of research belongs to the broader field of data assimilation, where available data and mathematical models are used both to estimate unknown parameters of a PDE (parameter estimation) or to correct model results (state estimation) [29, 30, 31]. In the latter case, as the problem considered in this work, the goal is to estimate time- and space-resolved fields over the whole computational domain or on a subset of interest.
For the forward modeling, we use a third-order weighted essentially non-oscillatory finite volume method (WENO3) combined with an optimized second-order backward differentiation formulation (BDF2-opt) with an implicit scheme to handle the temporal evolution. This approach ensures both accuracy and stability. Efficient implementations of these methods and high-performance computing techniques allow accurate and fast approximation of thermally-coupled flows in complex geometries. However, the need to systematically repeat these simulations, a key stage for the inverse problem, calls for the usage of Reduced-Order Modeling (ROM) strategies to reliably decrease the dimensionality of the problem, preserving its fundamental characteristics, using techniques such as the reduced-basis method [32], or proper orthogonal decompositions [33] (see e.g. [34, 35], for a review on ROM strategies).
As an inverse technique, we employ a regularized least-squares approach based on the so-called Parametrized Background Data Weak (PBDW) method, see, e.g., [36, 37, 38, 39]. In this approach, the optimization problem to be solved for the state estimation is based on the projection of the physics on a reduced-dimensional space that must reproduce the observed data in a weak form. Application of the PBDW in biomedical problems have been recently presented in [40, 41, 42, 43]. In this work, we use a methodology inspired by the PBDW to assimilate the temperature fields in the liquid-solid phase-change scenario, including turbulent flow conditions. In particular, we employ the ROM-based component of PBDW, which is equivalent to the full PBDW approach with infinite model regularization, as explained in [38], assuming no bias on the physical model. Additionally, we combine the state estimation with a Greedy algorithm for the optimal placement of temperature sensors. Greedy algorithms have been used for experimental design in the past years. In [44], a Bayesian design is used for seismic full waveform inversion. In [45], a greedy algorithm was employed using a minimal variance error for parameter estimation. In [46], optimal sensors are computed targeting directly a mean squared error between the model and measurements. Greedy methods have been combined with the PBDW in, e.g., [47, 48, 38]. In contrast with these works, our proposal solely relies on available a priori results for the regularized least-squares reconstruction, prioritizing sensor locations that maximize the observability of the reduced-order model.
In summary, this article proposes a temperature estimation framework from optimally placed synthetic data and a regularized least-squares estimator through a physics-informed projection, using turbulent simulations of the airflow surrounding the freezing food. The temperature estimation algorithm is proven valid for extrapolating fields in non-observable regions, keeping good reconstruction errors due to a properly designed sensor placement strategy. As a further application of the fully reconstructed field, we also show the capabilities of the algorithm to predict temperature-related quantities, such as Nusselt numbers, freezing time, and convective coefficients, aspects of high interest in food engineering. This proof-of-concept article will specifically use synthetically generated measurements from a thermal camera.
The rest of this manuscript is organized as follows. Section 2 states the problem, including the physical situation studied in this work, the governing equations solved and the models used, and the initial and boundary conditions. Section 3 summarizes the finite volume method used as a direct strategy, focusing on the distinctive aspects with respect to standard finite volume solvers used in the literature for food process modeling, and a standard validation of the method with well-established numerical and experimental benchmarks. In Section 4, an inverse pipeline for state estimation is presented, including the reduced-order strategy used and a fully devoted sub-section where an optimal sensor placement strategy is adapted for the problem at hand. In Section 5, the feasibility of the proposed inverse method to estimate states in processes involving food freezing is evaluated exhaustively using different sensor positions and highlighting the representativeness of the methodology for quantities of interest to the food industry. We finally draw concluding remarks in Section 6.
2 Mathematical modeling of the conjugate food freezing process
2.1 Model setting
We model the freezing process of food inside a convective freezer that circulates cold air between the cabinet and the refrigeration unit through inlet and outlet ducts. The conjugate dynamics of the process result from the simultaneous interplay of forced heat convection in the turbulent airflow and the nonlinear heat transfer coupled with the liquid-to-solid phase change process within the food slab. This setup is illustrated in Figure 1.
The freezer is assumed to have a cabinet of width $W=0.35$ m, height $H=0.4$ m, and depth $D=0.45$ m. Cold air coming from the refrigeration unit is forced into the cabinet by a fan at a constant velocity $U_{in}$ and temperature $T_{C}$ through an inlet duct of height $H_{in}=0.12$ m, and it leaves the cabinet through an outlet duct of height $H_{out}=0.06$ m. An idealized hexahedral food slab of width $W_{f}=0.08$ m, height $H_{f}=0.03$ m, and depth $D_{f}=0.08$ m, is placed on a glass shelf of width $W_{s}=0.25$ m and thickness $d_{s}=0.002$ m, located at the mid-height of the cabinet. Also, we assume that the walls of the cabinet exchange heat with the environment by conduction through the insulated layers and by natural convection with the surrounding air. The rear wall that separates the cabinet from the refrigeration unit is considered adiabatic.
Figure 1: Left: Sketch of the physical domain for the food freezing process. Right: Details of the two-dimensional section considered in this study.
In the present work, we focus on using an inverse pipeline to recover temperature fields from low-quality measurements that require building a dataset with multiple direct simulations. Therefore, we only consider a two-dimensional setup at the mid-section of the freezer (Figure 1, right). This two-dimensional model has been proven to represent with sufficient accuracy the food freezing process by turbulent forced and natural convection of airflow, see, e.g., the specialized references [10, 49]. In what follows, we will denote with $\Omega_{f}\subset\mathbb{R}^{2}$ and $\Omega_{s}\subset\mathbb{R}^{2}$ the fluid-flow and food domains, respectively, and $[0,t_{f}]$ is the physical time interval considered for the simulation.
2.2 Airflow dynamics
The airflow in the fluid domain is assumed to be described by the incompressible unsteady Reynolds-averaged Navier–Stokes equations (URANS) for velocity $\boldsymbol{u}$ and pressure $p$, coupled to a heat convection equation for the air temperature $T$. Using a Boussinesq approximation for the buoyancy forces to deal with the thermal coupling, the airflow obeys the following governing equations:
$\nabla\cdot\boldsymbol{u}=0\qquad\text{in}\ \Omega_{f}\times(0,t_{f}],$ (1)
$\rho\partial_{t}\boldsymbol{u}+\rho\boldsymbol{u}\cdot\nabla\boldsymbol{u}-\nabla\cdot[2(\mu+\mu_{t})\nabla^{s}\boldsymbol{u}]+\nabla p=\boldsymbol{f}\qquad\text{in}\ \Omega_{f}\times(0,t_{f}],$
$\rho\partial_{t}T+\rho\boldsymbol{u}\cdot\nabla T-\nabla\cdot[(\lambda/C_{p}+\sigma_{T}\mu_{t})\nabla T]=0\qquad\text{in}\ \Omega_{f}\times(0,t_{f}],$
In Eq. (1) $\rho$, $\lambda$, $C_{p}$, $\mu$, and $\beta_{ref}$ denote the density, thermal conductivity, specific heat, and viscosity of the air, respectively. These physical properties are assumed as constant at a reference temperature of $T_{ref}=0$$\degree$°C. The source term in the momentum equation,
$\boldsymbol{f}=\rho\beta_{ref}\left(T-T_{ref}\right)\boldsymbol{g}+\frac{2}{3}\nabla\cdot(\rho k),$
includes the buoyancy and the turbulent source terms, where $\beta_{ref}$ is the air’s thermal expansion coefficient and $k$ is the turbulent kinetic energy. The additional turbulent viscosity $\mu_{t}$ introduced by the URANS method is described by the $k$-$\omega$ SST model as a function of the turbulent kinetic energy and specific turbulence dissipation rate ($\omega$) as
$\mu_{t}=\rho\,\frac{a^{*}k}{\max(a_{1}\omega,SF_{2})}\,,$ (2)
while two additional scalar partial differential equations are introduced to close the model:
$\rho\partial_{t}k+\rho\boldsymbol{u}\cdot\nabla k-\nabla\cdot(\mu+\sigma_{k}\mu_{t})\nabla k=\tilde{P}_{k}-\beta^{*}\rho k\omega\qquad\text{in}\ \Omega_{f}\in(0,t_{f}]$ (3)
$\rho\partial_{t}\omega+\rho\boldsymbol{u}\cdot\nabla\omega-\nabla\cdot(\mu+\sigma_{\omega}\mu_{t})\nabla\omega=\frac{\rho\gamma}{\mu_{t}}\tilde{P}_{k}-\rho\beta\omega^{2}+C_{k\omega}\qquad\text{in}\ \Omega_{f}\in(0,t_{f}].$
In Eq. (2), $S=\sqrt{2\nabla^{s}\boldsymbol{u}:\nabla^{s}\boldsymbol{u}}$, $a^{*}=0.31$, and $F_{2}$ is an adjustable function of the model. In Eq. (3), $\beta^{*}=0.09$, $\tilde{P}_{k}=\min(\tau:\nabla\boldsymbol{u},10\beta^{*}\rho\omega k)$, $C_{k\omega}=2\sigma_{\omega,2}\omega^{-1}(1-F_{1})(\nabla k\cdot\nabla\omega)$ and $\sigma_{T}$, $\sigma_{k}$, $\sigma_{\omega}$, $\beta$, and $\gamma$ are functions varying as a convex combination of inner and outer constant values (close to the high gradient regions and far from them) with the blending function $F_{1}$, e.g., $\sigma_{\cdot}=F_{1}\sigma_{\cdot,1}+(1-F_{1})\sigma_{\cdot,2}$, where the indices $1$ and $2$ denote the inner and outer values, respectively. For the detailed model, we refer the reader to [50].
2.3 Food freezing process
The heat transfer and the liquid-solid phase change process in the food slab are described by a non-linear energy equation, assuming an incompressible and homogeneous material without considering mass transfer, such that:
$\partial_{t}(\rho_{s}C_{s}T_{s})-\nabla\cdot(\lambda_{s}\nabla T_{s})=0\qquad\text{in}\ \Omega_{s}\in(0,t_{f}].$ (4)
The internal convection of the liquid water contained in the muscle fiber and connective tissue of salmon can be neglected at the macro-level [20]. To account for the effect of latent heat during the phase change process in the salmon slab, the thermal properties (the density $\rho_{s}$, the specific heat capacity $C_{s}$, and the thermal conductivity $\lambda_{s}$) are considered as temperature-dependent, according to the effective heat capacity method [51, 52]. The temperature variability of the properties is depicted in Figure 2, assuming third-order fitting with (piecewise-constant) temperature-dependent coefficients:
$\psi=\sum_{j=0}^{3}a^{\psi}_{j}(T)T^{j},$ (5)
for $\psi=(\rho_{s}C_{s}$,$\lambda_{s}$). The values of the coefficients for the volumetric heat capacity $\rho_{s}C_{s}$ and thermal conductivity $\lambda_{s}$ are presented in Table 1.
Figure 2: Temperature dependency of the thermal properties of salmon meat. Left: density ($\rho_{s}$), center: specific heat capacity ($C_{s}$), right: thermal conductivity ($\lambda_{s}$).
Property $T\in$ [-25°C, -5°C] (-5°C, -3.5°C] (-3.5°C, 0°C] (0°C, 25°C]
$\rho_{s}C_{s}$ $a_{0}$ 4.298×$\times 10^{7}$ 3.046×$\times 10^{10}$4.604×$\times 10^{6}$ 4.365×$\times 10^{6}$
$a_{1}$ 7.166×$\times 10^{6}$ 1.906×$\times 10^{10}$1.037×$\times 10^{8}$3.228×$\times 10^{5}$
$a_{2}$ 4.199×$\times 10^{5}$ 4.002×$\times 10^{9}$1.095×$\times 10^{8}$ 2.535×$\times 10^{4}$
$a_{3}$ 8.031×$\times 10^{3}$ 2.817×$\times 10^{8}$4.260×$\times 10^{7}$5.838×$\times 10^{2}$
$\lambda_{s}$ $a_{0}$ 5.654×$\times 10^{-1}$3.276×$\times 10^{-1}$ 5.202×$\times 10^{-1}$ 5.122×$\times 10^{-1}$
$a_{1}$9.014×$\times 10^{-2}$5.611×$\times 10^{-1}$ 1.181×$\times 10^{-2}$3.900×$\times 10^{-3}$
$a_{2}$4.038×$\times 10^{-3}$9.160×$\times 10^{-2}$ 2.707×$\times 10^{-2}$ 3.585×$\times 10^{-4}$
$a_{3}$6.721×$\times 10^{-5}$5.767×$\times 10^{-3}$ 1.081×$\times 10^{-3}$7.767×$\times 10^{-6}$
Table 1: Coefficients for the polynomial functions of volumetric heat capacity ($\rho_{s}C_{s}$) and thermal conductivity ($\lambda_{s}$) of salmon for the different temperature ranges.
2.4 Boundary, initial, and interface conditions
Boundary conditions at the inlet $\partial\Omega_{in}$ are prescribed using constant air velocity $U_{in}$ and constant temperature $T_{C}$, which are assumptions based on: (i) the airflow enters horizontally with an average velocity that correlates with typical mass flow rates of medium-sized fans, and (ii) the refrigeration unit rapidly cools the air to the set freezing point during most of the charging process.
[…]
$\Omega$ Velocity Temperature Kinetic e. Dissip. rate of $k$
$\partial\Omega_{in}$: $\boldsymbol{u}=(U_{in},0)$ $T=T_{C}$ $k=(0.02U_{in})^{2}$ $\omega=\sqrt{k}/\beta^{*}H_{in}$
$\partial\Omega_{out}$: $\nabla\boldsymbol{u}\cdot\boldsymbol{n}=0$ $\nabla T\cdot\boldsymbol{n}=0$ $\nabla k\cdot\boldsymbol{n}=0$ $\nabla\omega\cdot\boldsymbol{n}=0$
$\partial\Omega$: $\boldsymbol{u}=0$ $\lambda\nabla T\cdot\boldsymbol{n}=(T_{ext}-T_{w})/R_{W}$ $k=0$ $\omega=800\mu/\rho d_{wall}^{2}$
$\partial\Omega_{sf}$: $\boldsymbol{u}=0$ $\lambda\nabla T\cdot\boldsymbol{n}=\lambda_{s}\nabla T_{s}\cdot\boldsymbol{n}$ $k=0$ $\omega=800\mu/\rho d_{wall}^{2}$
Table 2: Boundary conditions at the inlet duct $\partial\Omega_{in}$, the outlet duct $\partial\Omega_{out}$, the freezer walls $\partial\Omega$, and at the solid-fluid interface $\partial\Omega_{sf}$; $d_{wall}$ is the distance from the walls to the closest cell centroid.
Initial conditions in the fluid domain $\Omega_{f}$ are given by zero flow velocity, zero pressure field, and zero values of the turbulent variables, i.e. $\boldsymbol{u}_{0}=0$, $p_{0}=0$, $k_{0}=0$, and $\omega_{0}=0$. Also, the fluid and solid domains are in thermal equilibrium with the environment, i.e., $T_{0}={T_{s}}_{0}=T_{ext}$.
The inverse pipeline of the present work requires a dataset of snapshots from a set of time-dependent direct simulations with a variable space of parameters. Therefore, the magnitudes of the main boundary conditions are varied within a fixed range: $U_{in}\in[0.15,0.25]$ m-1, $T_{C}\in[-18,-26]\ ^{\circ}\text{C}$, $T_{ext}\in[18,26]\ ^{\circ}\text{C}$, and $h_{ext}\in[0.4,1.2]\ \text{W}\text{m}^{-2}\text{K}^{-1}$. The detailed procedure to construct this five-dimensional set is explained in Section 5.
3 Direct problem
3.1 Finite volume method
To compute a numerical solution, the system of equations (1)-(3), and (4) is discretized in space using the finite volume method (FVM) by decomposing the computational domain into $N_{v}$ non-overlapping cells $\Omega_{i}$. Integrating the PDEs over each cell and applying the divergence theorem yields:
$\int_{\Omega_{i}}\left(\rho\partial_{t}\phi\right)\mathrm{d}\Omega_{i}=\int_{S_{i}}\left(\rho\boldsymbol{u}\phi\right)\cdot\boldsymbol{n}\,\mathrm{d}S_{i}+\int_{S_{i}}\left(\Gamma\nabla\phi\right)\cdot\boldsymbol{n}\,\mathrm{d}S_{i}+\int_{\Omega_{i}}\boldsymbol{f}\,\mathrm{d}\Omega_{i},$ (6)
where $\phi$ stands for any state variable ($\boldsymbol{u},T,k,\omega$), $\Gamma$ is the corresponding diffusion coefficient, $\boldsymbol{f}$ stands for the corresponding forcing term, $S_{i}$ is the cell boundary, and $\boldsymbol{n}$ is the outward normal unit vector to $S_{i}$. We use a non-uniform structured mesh with a staggered grid approach that stores the cell-averaged value of scalars at the cell centroid and velocities at the cell faces.
The cell face reconstruction is based on a third-order weighted essentially non-oscillatory method (WENO3) [55] that uses two adaptive weights, $w_{0},w_{1}\geq 0$, with $w_{0}+w_{1}=1$, and two-point sub-stencils, $S_{i}^{0}(\phi_{i-1},\phi_{i})$ and $S_{i}^{1}(\phi_{i},\phi_{i+1})$, such that: $\phi_{S_{i}}=w_{0}S_{i}^{0}+w_{1}S_{i}^{1}$. Each sub-stencil approximates $\phi_{S_{i}}$ using a linear interpolation through the respective stencil points (cell centroids) in the non-uniform mesh. The weights are computed as a function of the state variable as:
$w_{k}=\frac{\alpha_{k}}{\alpha_{0}+\alpha_{1}}\ ;\ \alpha_{k}=\frac{d_{k}}{\left(\beta_{k}+\epsilon\right)^{2}}\ ;\ \beta_{k}=\left(\phi_{i-1+k}-\phi_{i+k}\right)^{2},\qquad\text{for}\ k=0\ \text{and}\ 1,$
where $d_{1}=\frac{1}{3}$, $d_{2}=\frac{2}{3}$, and $\epsilon=10^{-6}$. Finally, we use the Lax–Friedrichs procedure to ensure a consistent monotone physical flux at the cell faces, and we use a second-order central difference method for the gradient reconstruction at cell faces.
The time discretization is based on a uniform time step dividing the considered time interval $\left[0,t_{f}\right]$ into $N_{t}$ sub-intervals of size $\mathrm{\delta}t=t^{n+1}-t^{n}$. The state variables on the right-hand side of (6) are treated implicitly, and the time derivative is approximated using an optimized backward differentiation formula (BDF2-opt). This method is unconditionally stable and uses four previous time-levels, such that:
$\partial_{t}\phi=\delta t^{-1}\left(c_{1}\phi^{n+1}+c_{2}\phi^{n}+c_{3}\phi^{n-1}+c_{4}\phi^{n-2}\right),$
where the coefficients are calculated by a linear combination of a classical second-order and a third-order BDF schemes [56, 57], such that: $c_{i}=\chi c_{i}^{\textrm{BDF2}}+(1-\chi)c_{i}^{\textrm{BDF3}}$. In all simulations considered in this, we set $\chi=0.52$.
Velocity and pressure are coupled using the segregated pressure-correction method SIMPLEC, and the set of resulting linearized algebraic equations is solved by the Bi-Conjugated Stabilized Gradient method (Bi-CGSTAB) with the ILU preconditioner. The maximum relative difference of all state variables between two successive iterations is used as converge criteria. The solution converges when that error is less than $10^{-6}$.
3.2 Verification of the finite volume solver
To verify the numerical method, we solve two well-established benchmark problems:
  • (P1)
    Freezing process of beef-meat inside a conventional freezer cabinet by turbulent natural convection of air, with a Rayleigh number of $Ra=6.81\times 10^{7}$ [21], and
  • (P2)
    Turbulent natural convection of air in a square cavity differentially heated on the sides, with $Ra=1.58\times 10^{9}$ [58].
Figure 3: Numerical results for the present direct solver (continuous line), for a previous direct solver [10] (dashed line), and experimental data (squares and triangles) for the two benchmark problems: (a) P1: Temperature evolution at two selected points inside the food, as illustrated in lower-left corner. (b) P2: Profiles of vertical velocity and Nusselt number at the mid-height and at the hot wall of the cavity, respectively.
Figure 3 $a)$ shows, for P1, the numerical results for the temperature at two control points in the food domain ($A$ located at the geometrical center and $B$ located at a quarter height from the lower surface) comparing the numerical results obtained with the method described in Section 3 with previous results from the authors [10] and with available experimental temperature measurements. Concerning the results reported in [10], the present method uses higher order schemes for the cell-face reconstruction and for the time integration (WENO3 and BDF2-opt) and includes a $k$-$\omega$ SST model that significantly improves the accuracy of results, reducing the RMS error from 3.88 °C to 2.41 °C at $T_{A}$, and from 4.27 °C to 1.86 °C at $T_{B}$.
[…]
3.3 Forward numerical simulations
This Section describes the results of a representative parameter set defined as the mean values of the range for each boundary condition, i.e., $U_{in}=0.2$ m/s, $T_{C}=-22\ ^{\circ}\text{C}$, $T_{ext}=22\ ^{\circ}\text{C}$, and $h_{ext}=0.8\ \text{W}\text{m}^{-2}\text{K}^{-1}$. We discuss the main features of the direct numerical results, focusing, in particular, on the challenges related to the forced heat convection within a turbulent airflow and to the liquid-to-solid phase change process of the food.
Figure 4 illustrates snapshots of the temperature, air velocity as $U=\sqrt{u_{1}^{2}+u_{2}^{2}}$, and turbulent kinetic energy fields at characteristic time instants: $t=0.001$ h (hours), $t=1.5$ h, $t=4.5$ h, and $t=8$ h. The results indicate that the inlet airflow causes rapid cooling of the air domain within the first minute, with some high-temperature stratifying isotherms in the upper region produced by natural convection effects. Contrary, the food cools down slowly due to its high thermal inertial caused by the latent heat during phase change. Moreover, both air velocity and turbulent kinetic energy fields indicate that at the very beginning the flow is more unsteady and chaotic, evidencing that the fluctuations and gradients of air velocity are notably higher. However, the flow stabilizes rapidly within the first hour and reaches a semi-steady state.
To understand the different time scales of the fluid dynamics and heat transfer processes, Figure 5 depicts the variation in time of volume-average and maximum values of temperatures (air and food), of food liquid fraction ($F_{pc}$), and of time-averaged and instantaneous fluctuations of velocity, i.e. $U$ and $U^{\prime}=\sqrt{(2/3)k}$. Analyzing the figure one sees the abrupt drop in air temperatures. In contrast, the average and maximum temperature in the food illustrates a characteristic freezing curve where the effect of the latent heat in the temperature-slope is clearly defined in the freezing stage. Regarding the evolution of volume-average and maximum values the two components of the air velocity in a URANS model, the most dynamic patterns of $U$ and $U^{\prime}$ are observed at the first seconds of the freezing process, but then these values quickly stabilize until they reach a steady behavior over time. Moreover, the elevated value of $U_{max}$ is about four times higher than its volume-average value, demonstrating the presence of high-velocity gradients in the flow domain. Lastly, the relation between the instantaneous fluctuation and the time-average velocity, or turbulent intensity $I_{t}=U^{\prime}/U$, reaches volume-average and maximum values of 22% and 38%, respectively, indicating the convective-dominant nature of the problem. These points evidence the need for a suitable turbulent model that precisely computes the augmented diffusion of momentum and heat due to the turbulent mixing phenomenon.
Figure 4: Depictions of different snapshots of the temperature fields $(T)$, air velocity fields $(U)$, and turbulent kinetic energy fields $(k)$ for the representative set of parameters.
Figure 5: Left: Evolution of volume-averaged and maximum values of temperature in the air and food, and of the food liquid fraction. Right: evolution of volume-averaged and maximum values of the time-averaged air velocity $U$ and of the instantaneous fluctuation of velocity $U^{\prime}$.
Figure 6 (left) shows snapshots of the temperature and liquid fraction fields within the food at characteristic time instants: $t=0.2$ h (hours), $t=1$ h, $t=3$ h, and $t=5$ h. As expected, the liquid fraction field evolution is directly correlated with the isotherms since it is a temperature-derived quantity. Another relevant aspect observed in the results is that the freezing center moves during the freezing process. Initially, the high-temperature isotherms are concentrated just below the geometric center, and during the phase change, they are rather concentrated close to the right surface of the food. This behavior occurs because the cold air flow slips along the upper wall and hits the food on the left, while the lower right corner is located precisely in a flow recirculation region, which weakens heat transfer.
To highlight the spatial differences in the dynamics of the food freezing process, Figure 6 (right) depicts the temporal evolution of temperature, liquid fraction, and volumetric heat capacity at two fixed points within the food domain. These points are conveniently selected as the freezing center ($P1$), and in the upper corner region ($P2$) where the freezing rate is expected to be highest. The difference in the path of temperature and liquid fraction evolution between the two points within the food shows the relevance of the locality in the phase change phenomenon. The center region experiences a more isothermal process than in the edges of the food, indicating a lower heat transfer rate in that area. Consequently, the maximum values of volumetric heat capacity related to the latent heat occur more slowly and at different periods of time in the freezing process.
Figure 6: Left: Snapshots of temperature and liquid fraction fields within the food domain. Right: local evolution of the temperature, liquid fraction, and volumetric heat capacity at two control points.
4 Inverse problem
We aim to reconstruct the temperature field from the available data by solving an optimization problem that minimizes the discrepancy between the model and the measurements. Specifically, the problem is formulated as follows: Given an experimental dataset $\ell\in\mathbb{R}^{m}$, in a fixed time, where $m$ is the number of measurements (pixels in a picture, for instance), find a field $T^{*}$ (e.g., the reconstructed temperature) such that:
$T^{*}=\operatorname*{arg\,min}_{T\in\mathbb{R}^{\mathcal{N}}}\|\ell-\boldsymbol{W}^{T}T\|_{2}.$ (7)
The minimizer $T^{*}$ is sought in an $\mathcal{N}$-dimensional space, where $\mathcal{N}$ is the number of degrees of freedom from the numerical solution considered for the forward problem (proportional to the number of cells $N_{v}$ in the finite volume discretization), and where $\boldsymbol{W}\in\mathbb{R}^{\mathcal{N}\times m}$ is a suitable matrix representing the observation process, herein assumed linear on the state variable $T$.
The main proposal of this work is a procedure for estimating temperature fields from experimental data. In this context, there are several alternatives for handling the input data, which we can summarize in two types of measurements: Point-wise measurements and averaged temperature sensors in low-resolution temperature fields from thermo-cameras. In any case, the sensor gives local averages of the temperature field to be estimated, and therefore, the linearity assumption on the state to be estimated is reasonable.
As schematically depicted in Figure 7, measuring a high dimensional temperature state is understood as applying the observer matrix $\boldsymbol{W}$ on a temperature state. The matrix columns, otherwise understood as the Riesz representers of the measurements [42], are indicatrix functions on the data pixels, i.e., for every thermo-camera pixel $\Omega_{i}\subset\Omega$ we have:
$\ell_{i}(\Omega_{i};x)=\left\{\begin{aligned} 1/\|\ell_{i}(\Omega_{i})\|,&\quad x\in\Omega_{i},\\ 0,&\quad x\in\Omega\backslash\Omega_{i},\end{aligned}\right.$
for $i=1,\ldots,m$ pixel measurements.
Figure 7: The observation model as a linear operation on the target state. Local averages of the temperature are computed at each pixel of the thermo-camera data $\Omega_{i}\subset\Omega$.
4.1 Least-squares reconstruction with ROM regularization
Typically, the number of available measurements is smaller than the dimension of the numerical solution ($\mathcal{N}>m$). As a consequence, problem (7) does not have, in general, a unique solution. This can be seen by developing the optimal conditions of the optimization problem and noticing that the resulting linear system, i.e.,
$\boldsymbol{W}\boldsymbol{W}^{T}T=\boldsymbol{W}\ell,$ (8)
is not invertible, as the projection matrix $\boldsymbol{W}\boldsymbol{W}^{T}$ has at most $m$ non-zero eigenvalues. Otherwise stated, for the usual case $\mathcal{N}>m$, there are infinitely many temperature fields whose measurements are equal. To overcome this issue, the strategy adopted in this work relies on seeking the minimizer in a suitable subspace of dimension $n\ll\mathcal{N}$ (ROM - reduced-order model) defined by the basis:
$\boldsymbol{\Phi}=\left(\phi_{1}|\ldots|\phi_{n}\right)\in\mathbb{R}^{\mathcal{N}\times n}\,.$
Hence, the minimization problem (7) is replaced by:
$c^{*}=\operatorname*{arg\,min}_{c\in\mathbb{R}^{n}}\|\ell-\boldsymbol{W}^{T}\boldsymbol{\Phi}c\|_{2}.$ (9)
The optimality conditions for the reduced-order problem (9) yield the well-known normal equations [42]:
$\boldsymbol{G}^{T}\boldsymbol{G}c=\boldsymbol{G}^{T}\ell,$ (10)
where $\boldsymbol{G}:=\boldsymbol{W}^{T}\boldsymbol{\Phi}$ is a cross-Gramian matrix. Then, the temperature field in the original space can be recovered doing $T^{*}=\boldsymbol{\Phi}c^{*}$.
Properly built, the reduced-order space defined by $\boldsymbol{\Phi}$ will maintain the most relevant features of the original transport problem described by Eq. (1). A practical approach to defining a suitable physics-based basis $\boldsymbol{\Phi}$ will be described in more detail in Section 4.2.
The dimensional reduction has two important implications:
  • Speed-up: The reduced model in $\boldsymbol{\Phi}$ has a considerably smaller dimension than the original, full-order problem. Provided that this low-dimensional space can represent sufficiently well the physics of the original problem, the overall complexity is considerably reduced, since only an $n\times n$ system shall be inverted.
  • Physics-driven recovery: The reduced-order model is tightly coupled to the original physics. Hence, it regularizes the otherwise non-invertible problem (7) by reducing the number of unknowns with a physics-informed search space for the optimization problem.
The analogous of problem (7) in a continuous and infinite-dimensional setting has been studied, from a general point of view, in [36, 59, 60]. The regularized least-squares approach used in this article is one of the numerical ingredients of the Parametrized Background Data-Weak (PBDW) method proposed in [36], or it can be seen as the perfect-model (without bias) case explained in [38]. The analysis of the resulting low-dimensional problem reveals that the condition $\mathcal{N}>m>n$—indicating that the number of observations exceeds the dimension of the reduced-order model—is necessary to ensure well-posedness. This condition will be assumed valid throughout the remainder of this article.
Thus, the reconstruction problem is considered to have two stages. An offline (or training) stage dedicated to the simulation of several direct problems, plus the computation of the ROM and the observation matrix $\boldsymbol{W}$. In the offline stage, most of the computational cost is focused, in contrast with the online phase, where the sole inversion of a $n\times n$ system of equations enables real-time computations.
4.2 Reduced-order modeling
This section discusses the generation of the reduced-order model $\boldsymbol{\Phi}$ using a parametrized singular value decomposition (SVD) [61] of a set of snapshots from the forward problem (named in certain contexts PCA – principal component analysis – or POD – proper orthogonal decomposition – in its generalization to arbitrary infinite-dimensional functional settings). The goal is to define a low-dimensional space that can sufficiently well approximate the physics of the phase change problem and the turbulent heat convection developed by the airflow. For this purpose, the first step consists of concatenating snapshots
$\boldsymbol{A}=\left(T(\theta_{1})|\ldots|T(\theta_{k})\right),$
of the numerical solution of Eq. (1) for a (large) amount of configurations of the parameters defining the underlying model $\theta_{1},\ldots,\theta_{K}$. In the above notation, the vector $\theta_{i}\in\mathbb{R}^{p}$ ($p\in\mathbb{N}$) defines a generic set of parameter values for (1), e.g., time, fan configurations, boundary conditions, phase-change parameters, food-specific coefficients, among others.
The basis of the reduced-order space $\boldsymbol{\Phi}$ is then computed using the unique and truncated singular value decomposition of the snapshot matrix $\boldsymbol{A}$, namely:
$\boldsymbol{A}=\boldsymbol{\Phi}\Sigma\boldsymbol{V}^{T}=\underbrace{\begin{bmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&\cdots&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&\cdots&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}\\ \vdots&\vdots&\ddots&\vdots\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&\cdots&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&\cdots&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}\end{bmatrix}}_{\boldsymbol{\Phi}\ (\mathcal{N}\times n)}\underbrace{\begin{bmatrix}\sigma_{1}&0&\cdots&0\\ 0&\sigma_{2}&\cdots&0\\ \vdots&\vdots&\ddots&0\end{bmatrix}}_{\Sigma\ (n\times n)}\underbrace{\begin{bmatrix}{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}&\cdots&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}\\ {\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}&\cdots&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}\\ \vdots&\vdots&\ddots&\vdots&\vdots\end{bmatrix}^{T}}_{\boldsymbol{V}^{T}\ (n\times k)}$ (11)
defining $\boldsymbol{\Phi}$ using the first $n$ columns of the full decomposition. We will call $\sigma_{i}$ the diagonal entry $\Sigma_{ii}$ of the matrix $\Sigma$. We see that the core idea of the SVD regularization for the inverse technique relies upon constraining our optimization search to a linear combination of the columns in $\boldsymbol{\Phi}$ (red squares in (11)).
4.3 Greedy algorithm for optimal sensor placement
As shown in, e.g., [40, 36], the following a-priori error bound for the reconstructed temperature field $T^{*}$ holds
$\frac{\|T-T^{*}\|}{T}\leq\left(\inf_{c\in\mathbb{R}^{n}}\frac{\|\boldsymbol{G}c\|}{\|\boldsymbol{\Phi}c\|}\right)^{-1}\frac{\|A-A_{n}\|_{\text{FRO}}}{\|A\|_{\text{FRO}}}$ (12)
where $\boldsymbol{A}_{n}$ is the $n$-th low rank approximation of $\boldsymbol{A}$ and ${\|\cdot\|_{\text{FRO}}}$ denotes the Frobenius norm [61]:
${\|A\|^{2}_{\text{FRO}}}=\mathrm{tr}(A^{T}A)=\sum_{i,j}A_{ij}^{2}\,.$
The bound (12) shows where the condition $m>n$ comes from, indicating that a reduced model that is too enriched would lead to non-observable modes, a critical issue for well-posedness in data assimilation problems. Using standard arguments, the second term in the right-hand-side of (12) can be expressed in terms of the singular values of $\boldsymbol{A}$:
$\frac{\|A-A_{n}\|_{\text{FRO}}}{\|A\|_{\text{FRO}}}=\frac{\sqrt{\sum_{i=n+1}^{\text{max}\{\mathcal{N},K\}}\sigma_{i}^{2}}}{\sqrt{\sum_{i=1}^{\text{max}\{\mathcal{N},K\}}\sigma_{i}^{2}}}$
For the first term, let us consider the spectral induced $\ell^{2}$ norm, considering the singular value decomposition $\boldsymbol{G}=\hat{U}\hat{S}\hat{V}^{T}$, and as usual, denote with $\hat{S}_{i}$ the entry $\hat{S}_{ii}$ of the matrix $\hat{S}$. By construction, the columns of $\hat{V}$ are a basis of $\mathbb{R}^{n}$, allowing us to write, for any $c\in\mathbb{R}^{n}$, $c=Vd$ for a unique $d\in\mathbb{R}^{n}$. Since both the columns of $V$ and of $\boldsymbol{\Phi}$ are orthonormal, we obtain
$\left(\frac{\|\boldsymbol{G}c\|}{\|\boldsymbol{\Phi}c\|}\right)^{2}$ $=\frac{d^{T}\hat{V}^{T}\boldsymbol{G}^{T}\boldsymbol{G}\hat{V}d}{d^{T}V^{T}\boldsymbol{\Phi}^{T}\boldsymbol{\Phi}Vd}=\frac{d^{T}\hat{S}^{2}d}{d^{T}d}=\frac{\sum_{i=1}^{n}d_{i}d_{j}\hat{S}_{i}^{2}}{\sum_{i=1}^{n}d_{i}^{2}}$ (13)
$\geq\frac{\hat{S}_{n}^{2}\sum_{i=1}^{n}d_{i}^{2}}{\sum_{i=1}^{n}d_{i}^{2}}=\hat{S}_{n}^{2},\;\forall c\in\mathbb{R}^{n},$
where we have also used the fact that the singular values are decreasing and positive. The inequality (13) allows to conclude that
$\inf_{c\in\mathbb{R}^{n}}\frac{\|\boldsymbol{G}c\|}{\|\boldsymbol{\Phi}c\|}\geq\hat{S}_{n},$
and using (12) to obtain the following a priori bound for a given dimension $n$ of the reduced-order space
$e(n)=\hat{S}_{n}^{-1}\frac{\sqrt{\sum_{i=n+1}^{\text{max}\{\mathcal{N},K\}}\sigma_{i}^{2}}}{\sqrt{\sum_{i=1}^{\text{max}\{\mathcal{N},K\}}\sigma_{i}^{2}}}.$ (14)
Equation (14) has two main implications for the properties of the temperature reconstruction. On the one hand, better approximation properties of the reduced-order model lower the enumerator of the error bound. On the other hand, the term $\hat{S}_{n}^{-1}$ (the last singular value of the cross-Gramian matrix $\boldsymbol{G}$) indicates that increasing the dimension of the reduced-order space, i.e., decreasing $\hat{S}_{n}^{-1}$, lowers the quality of the expected result. This trade-off can be interpreted as a lack of observability of the reduced-order space if the dimension is too large. From the point of view of linear algebra, the observability can be defined as the angle between the subspace spanned by the columns of the matrix $\boldsymbol{W}$ and the reduced-order model spanned by the columns of $\boldsymbol{\Phi}$ [59].
We will use (14) as a criterion to improve the ROM observability and to compute the optimal placement of measurement sensors. In the rest of this section, we describe an iterative approach to systematically increase the dimension of the matrix $\boldsymbol{W}$ (i.e., the number of measurements) by locating the sensors in such a way as to maximize the observability (i.e., minimizing (14)) for a fixed dimension $n$ of the reduced-order model.
Using a criteria based on (14) to select the optimal dimension of the ROM and data assimilation has been already applied in other contexts [40, 41, 42]. In [47], a classical QR pivoting strategy is employed (thus limiting the algorithm to square systems of equations), whereas in [38], a criteria based on (14) is used, nonetheless with certain differences: First, with our method, we do select the optimal measurements based on a fixed dimension for the ROM (which could be necessary for certain engineering applications, where a certain level of approximation should be ensured for the reduced basis). Second, in [38], the measurement addition is based upon optimizing the least-stable mode iteratively instead of using all modes at once, as we propose here. Both approaches may deliver similar results as $m$ increases, but for a fixed amount of measurements (as in many engineering scenarios), we compute the global stability using the full reduced basis, thus ensuring that the conditioning is optimal for fixed $n$ and $m$.
Then, for a given $n$, we aim at selecting the location of $m\geq n$ sensors from a pool of $M$ possible choices over the observable region $\Omega$. The proposed approach follows the underlying idea of a Greedy method, adding the sensors progressively and looking at each stage for the location of a new sensor that yields the smallest a-priori bound (14). Let us recall that, as seen in 4, a measurement sensor can be seen as a linear operation on the temperature field. Thus, considering the Riesz representation of this linear functional, in what follows, we will refer to a sensor as a column of the matrix $\boldsymbol{W}$. Moreover, we will denote as $\mathcal{S}=\{\ell_{1},\ldots,\ell_{M}\}$ a set of all possible sensor locations. The step-by-step process is shown in the blue box (Algorithm 1). Besides optimizing the conditioning of the matrix in the normal equations, this approach for the selection of additional modes has the physical meaning of adding experimental measurements at locations where the most relevant dynamics can be observed.
Algorithm 1: Greedy Sensor Placement Input: ROM basis $\boldsymbol{\Phi}$ (and its fixed dimension $n$). Pool of feasible sensor positions $\mathcal{S}$. Step 1 Select the first sensor as: ^$\hat{\ell}_{1}:=\operatorname*{arg\,max}_{w\in\mathcal{S}}\|w^{T}\boldsymbol{\Phi}\|.$ Step 2 Select the next $n-1$ sensors (necessary to ensure well-posedness) as: ^$\hat{\ell}_{i}:=\operatorname*{arg\,max}_{w\in\mathcal{S}}\sigma_{i}(\hat{\mathbf{W}}(w)^{T}\boldsymbol{\Phi}),\;\text{for }i=2,\dots,n,$ where $\hat{\mathbf{W}}(w):=\big{[}\hat{\ell}_{1}\mid\cdots\mid\hat{\ell}_{i-1}\mid w\big{]}$ is the $\mathcal{N}\times i$ observation matrix for a candidate $i$-th sensor $w$, and $\sigma_{i}(B)$ denotes the $i$-th singular value of any matrix $B$. Step 3 For $i>n$, locate any additional sensor maximizing the smallest non-zero singular value of the resulting sensor matrix, i.e., ^$\hat{\ell}_{i}:=\operatorname*{arg\,max}_{w\in\mathcal{S}}\sigma_{n}(\hat{\mathbf{W}}(w)^{T}\boldsymbol{\Phi}).$
5 Numerical examples
This section focuses on evaluating the proposed inverse methodology in various settings. First, we validate the reconstruction of the full temperature field from synthetic thermo-camera measurements acquired over the whole computational domain. Secondly, we show how the method can be used to extrapolate the temperature field in the salmon slice using only synthetic measurements within the airflow domain. Thirdly, we apply the proposed Greedy algorithm to determine optimal sensor placement, and verify its robustness as the number of measurements diminishes.
The pipeline of the reduced-order inverse estimation framework proposed in this work is illustrated in Figure 8. In the training stage, we construct a five-dimensional parameter set by gathering snapshots from 64 simulations of the direct model. The resulting time-dependent dataset is obtained by randomly varying the boundary conditions, i.e., $U_{in}$, $T_{C}$, $T_{ext}$, and $h_{ext}$, within a fixed range by using a uniform random distribution, as presented in the lower-left corner of Figure 8. The magnitude of these variation ranges is consistent with real operating conditions of convective freezers and was estimated from the literature review and technical information from previous works [49, 23]. Then, we train the reduced order space by using an SVD with 48 simulations from the previous dataset, resulting in a total of 1.2×$\times\dot{1}0^{4}$ snapshots. The first 15 modes of the resulting bases are depicted in Figure 9. Thereafter, we manufacture artificial low-quality measurements of the temperature fields from the direct model by degrading the resolution in sensors or pixels of size $2\times 2$ cm2 (see also Figure 7). Finally, using a Greedy algorithm, we find optimal sensor placements based on a-priori error optimization. In the reconstruction stage, we recover the time-dependent temperature fields from 16 synthetic measurements using direct simulations different from those within the training dataset, using the ROM-based least-squares approach described in the previous section.
For the computational implementation of the numerical experiments, for the finite volume scheme solution of the forward problem, we have used an in-house library called MVF2D [62]. For the full ROM and inverse pipeline, we have used the software MAD ([63, 64], chapter 5), used successfully for forward and inverse problems in previous works such as [42, 65, 66].
Figure 8: Flow diagram of the present inverse framework used to reconstruct the temperatures.
Figure 9: The first 15th ordered (from left to right and top to bottom) modes of the reduced-order model for the direct problem using 48 simulations and a singular value decomposition.
5.1 Reconstruction of temperature fields from thermocamera measurements
The first test considers full-domain synthetic measurements mimicking a thermocamera placed over the two-dimensional mid-plane of the freezer. This setup results in an image composed of $m=361$ pixels or sensors. Having this setup fixed for the measurement vectors $\ell_{i}$, one can evaluate the matrix conditioning of the normal equations (10) summarized in (14). The decay of the eigenvalues for the covariance matrix associated with the snapshot matrix is shown in Figure 10. The values and decay tendency of the eigenvalues are related to the ROM’s ability to approximate the original model’s physics. The rapid decay of the values suggests, in particular, that the SVD provides a good approximation of the temperature field in the phase change phenomenon, also with a limited number of modes.
The indicator $e(n)$ from (14) for the fixed measurement setup is shown in Figure 10b. Since this quantity is related to an upper bound for the average error when computing the reconstruction, we infer an optimal dimension for the reduced-order model of $n=111$. The curve also clearly shows how adding more modes affects the matrix conditioning, lowering the quality of the expected result due to the reduced observability of the ROM when $n$ approaches $m$.
(a) Eigen-values decay for $\boldsymbol{A}^{T}\boldsymbol{A}$
(b) A-priori error (14)
Figure 10: Reconstruction error study for 2D food-freezing numerical example. Optimal results are achieved according to this bound for $n=111$.
A comparison of the temperature fields at three different snapshots is provided in Figure 12, showing the ground truth $T_{\text{GT}}$ (not included in the reduced-order model training set), the measurements $\ell=\boldsymbol{W}^{T}T_{\text{GT}}$, and the reconstructions $T^{*}$. As a benchmark criteria, we introduce the relative $\ell^{2}$ error:
Relative 2-norm error (%)=$\text{Relative $\ell^{2}$-norm error (\%)}=\frac{\|T_{\text{GT}}-T^{*}\|}{\|T_{\text{GT}}\|}\times 100.$ (15)
which we evaluate for every test case. The average result for the 16 simulations is depicted with respect to time in Figure 11, where we can conclude an overall good reconstruction quality during the whole freezing process, thus validating our framework at a proof-of-concept stage.
In addition, we do test exhaustively the reconstruction algorithm using a single, randomly chosen, parameter configuration, with the following values: $U_{in}=0.2\ ms^{-1}$, $T_{C}=-21.7\ \degree C$, $T_{ext}=21.3\ \degree C$, and $h_{ext}=0.88\ Wm^{-2}K^{-1}$.
Figure 11: Temporal evolution of the average reconstruction error for the full test dataset of 16 simulations out of the training set.
We observe a good qualitative agreement between the ground truth and the reconstructed field during the whole freezing process, especially within the food domain, which is of utmost interest for the targeted application. The field reconstruction indicates that the turbulent airflow recirculation through the cabinet is properly described. Moreover, the congruence in the concentration of freezing fronts within the food demonstrates that the phase change process is adequately represented. However, slight differences are noticeable in the upper region of the cabinet, where the airflow is more chaotic due to turbulent fluctuations and the stratification by buoyant forces.
Figure 12: Temperature field evolution of (i) the ground-truth from the direct model, (ii) the full-domain measurements, and (iii) the reconstruction procedure at three characteristic instants of time.
In the food freezing process, the local estimation of heat transfer quantities is of great relevance given that the final quality of frozen food depends mainly on the local freezing rate and the surface heat fluxes it receives. Therefore, in this Section we compare the evolution of two quantities of interest and their local relative error associated only with fixed points within the food for thermocamera measurements. Later, in Section 5.3 we compare the full-domain errors in the form of the l2-norm. The local errors are calculated relative to the time-average values of the quantity of interest $\Phi$ at the point $P$, defined as:
$\Phi_{rel,error}(\%)=\frac{|\Phi_{P}-\Phi^{*}_{P}|}{|\overline{\Phi}_{P}|}\times 100$ (16)
[…]
By analyzing the time-dependent relative errors of temperature and freezing rate (see Fig. 13, bottom), we find that the control point in the freezing center presents larger errors than in the food boundaries for the whole process. We must consider that the freezing center represents the point where the freezing fronts concentrate after passing through the entire food domain, so it accumulates all the numerical errors in space and time of the nonlinear food phase-change phenomenon. Therefore, it can be considered that this area is prone to greater numerical errors, which could explain the lower quality reconstruction. Moreover, the local errors increase mainly in two characteristic periods of the freezing process: first, at the very beginning of the initial cooling stage (before 0.1 h), the relative error reaches 1.65%, and then, at the end of the freezing stage (from 4 to 6 h), where the relative errors get to 1.35%. Finally, the relative errors are directly correlated with the freezing rate, which is expected given that the abrupt changes in temperature require an improved temporal discretization. Nevertheless, the relative errors can also be related to the chaotic initial cooling stage of the cabinet which causes slight discrepancies, as illustrated in Figure 12, and with the end of the phase-change process that accumulates the local error previously discussed.
(a) Local evolution of temperatures and their relative errors.
(b) Local evolution of freezing rates and their relative errors.
Figure 13: Comparison of local quantities at three fixed points inside the food between the ground-truth and the reconstructions using thermocamera measurements.
5.2 Food temperature reconstruction with extrapolated data from the airflow domain
[…]
Figure 14: Left: Snapshots of the truncated field of measurements, and of zoomed regions of the ground-truth and the reconstruction fields. Right: Local temperature evolution and its relative error at two fixed points of interest within the food domain.
To test the reconstruction quality of derived quantities in the food, Figure 15 compares the contours of liquid fraction $F_{pc}$ and of volumetric heat capacity $\rho_{s}C_{s}$ at three snapshots between the ground-truth and the inverse framework without measurements in the food domain. The first obvious observation is that the distribution of both derived quantities follows the same pattern described by the isotherms: at the first instant of time, the freezing fronts start to advance from the lateral and upper surfaces, but the latent heat contours do not yet cover large regions of the food; then, at $t=1.5$ h, the freezing fronts concentrate in the center with a slight inclination towards the right lateral surface, while the specific heat contours show clear peaks due to the latent heat; finally, at $t=4$ h, there remains a small area that is not yet frozen, which still concentrates a latent heat peak.
Regarding the reconstruction quality, the results generally indicate that the contours advance at a similar speed but with a slight phase shift throughout the process. In addition, a flatter distortion in the shape of the contours is observed, especially at $t=1.5$ h. Considering the long duration of the process ($t\approx 10$ h) and the nonlinear nature of the temperature and derived quantities evolution, the reconstruction procedure achieves excellent physical representation concerning the ground-truth from the direct model.
Figure 15: Comparison of contours of liquid fraction and volumetric heat capacity at at three snapshots between the ground-truth and the inverse model with extrapolated sensors from the airflow domain.
This result is very relevant, as it allows us to conclude that the food temperature fields can be reliably estimated without intervening in food, foreseeing this numerical tool as a powerful instrument to complement physical sensors in the food industry.
5.3 Temperature reconstruction with an optimal sensor placement algorithm
In this section, we test the optimal sensor placement algorithm proposed in section 4.3 using the same set of parameters for the boundary conditions used in Section 5.2. We compare the reconstruction output of this algorithm with a regularly chosen observation region, as depicted in Figure 16.a. Nested squared observation regions are taken, excluding the food region, leading to 4 sensors set-ups, namely with 56, 162, 266 and 352 (full domain excluding food) measurements. In contrast, a sparser sensor selection is made when using the 3-step algorithms, as shown in Figure 16. Interestingly, the method tends to privilege both the food surroundings and the upper part of the computational domain, corresponding to the region where the turbulent airflow interacts with the food and with the stratified air in the upper part of the cabinet.
(a) Regular placement
(b) Greedy placement (picked sensors in gray)
Figure 16: Optimal sensor placement comparison with regular squared measured regions. The Greedy algorithm prioritizes the ROM observability, thus enabling a better stability of the inverse problem.
To evaluate and compare the performance of the proposed Greedy algorithm against regular sensor placement, we compute the relative l2-norm error of temperature defined in Eq. (15). The time-dependent and time-averaged values of the relative l2-norm error between regular and optimal sensor placements with different amount of measurements are compared in Figure 17 a), and 17 b), respectively.
The results demonstrates the robustness of the Greedy algorithms also when decreasing the number of measurements. We observe how properly placing the sensors improves the quality of the reconstructions during the entire freezing process, achieving errors close to 1% over the whole time-dependent dynamics. Figure 17.a shows that the optimal placement of 162 sensors achieves a similar error (in $\ell^{2}$-norm) over time as the regular placement with almost twice as many sensors ($m$ = 352). Moreover, when comparing the time-averaged errors for a given number of sensors (Figure 17.b), the Greedy algorithm outperforms the regular placement, especially for a small number of measurements (such as for $m=52$), reaching similar time-average errors that those obtained with full-domain measurements (thermocamera).
(a) Time-dependent $\ell^{2}$-norm error.
(b) Time-averaged $\ell^{2}$-norm error.
Figure 17: Comparison of the reconstruction quality of the regular sensor placement and Greedy algorithm for optimal sensor placement. The Greedy algorithm allows for considerable reduction of the number of sensors without any significant error increase.
Similar to the previous sections, we evaluate the reconstruction quality for two different local quantities of interest. This time, we compare the ground-truth with the Greedy algorithm for the three set of sensors, i.e. $GA_{1}$ for $m=52$, $GA_{2}$ for $m=162$, and $GA_{3}$ for $m=266$, for the local evolution of temperature in Figure 18 a), and for the local freezing rate in Figure 18 b). Additionally, we compute the accumulative relative error over time by integrating in time the Equation 16.
From Figure 18 a) and b), we can highlight that the local temperature evolution and the freezing rate are well reproduced with the Greedy strategy using $GA_{1}$ and $GA_{3}$ in points P1 and P3. For point P2, $GA_{3}$ depicts a better reconstruction quality, particularly in the freezing rate during the phase-changing stage. Regarding the accumulative relative error over time, we can highlight that half of the total temperature error and two-thirds of the total freezing rate error accumulated at the end of the process comes from the initial chaotic stage, which is associated with the unsteady turbulent fluid dynamics and the rapid cooling of the food boundaries. At this stage, the temporal derivative of the reconstructed field is of relevance. Despite this, the method reconstructs the temperature field very well and even achieves a good definition of its temporal evolution, which is not trivial since the proposed algorithm does not incorporate any condition regarding the temporal derivative of the reconstructed field.
(a) Local evolution of temperatures and their accumulative relative errors.
(b) Local evolution of freezing rates and their accumulative relative errors.
Figure 18: Comparison of local quantities for three fixed points inside the food between the ground-truth and the reconstruction with the Greedy algorithm using three sets of measurements.
6 Conclusions
This study addresses the inverse estimation of temperature fields in food freezing processes, a critical factor in ensuring product quality in the food industry. Numerical simulations of thermally-coupled flow systems, including liquid-to-solid phase change and turbulence modeling, have been used to generate a synthetic dataset and to create a physics-informed reduced-dimensional space to regularize the optimization problem for the state estimation . The approach allowed to effectively reconstruct the time-dependent temperature distribution within the food matrix from sparse sensor data, even when observations inside the food were completely excluded from the available dataset. The optimal placement of temperature sensors was determined using a new greedy algorithm based on available a-priori estimates on the reconstruction, enhancing estimation accuracy and minimizing the number of sensors required. The novelty of using this approach is two-fold. Firstly, this strategy has not yet been employed in the context of food freezing process. Secondly, the underlying mechanism of the Greedy approach is based on a theoretical work on a-priori bounds (14), and, unlike more traditional sensor placement approaches, it does not directly relies on online phase information. In other words, the strategy only requires the ROM and the position of the sensors, but not the values of the measurements themselves, allowing for general conclusions independent of the data being assimilated at a given iteration. The Greedy approach has been successfully validated for defining sensor locations that maximized data representativeness, offering a cost-effective solution for complex temperature monitoring applications in food preservation.
Future work will explore further refinements in the reconstruction algorithm, including model-bias [38], and/or measurement-bias [43], control over the time derivative of the reconstructed field, as well as real-time estimation techniques to advance food safety protocols.
Acknowledgments
E. Castillo acknowledges the support provided by ANID-Chile throughout the project FONDECYT 1210156. D.R.Q. Pacheco acknowledges funding by the Federal Ministry of Education and Research (BMBF) and the Ministry of Culture and Science of the German State of North Rhine-Westphalia (MKW) under the Excellence Strategy of the Federal Government and the Länder. F. Galarce acknowledges the support provided by DI-VINCI PUCV funding 039.482/2024. Diego. Rivera acknowledges the support granted by the DICYT-POSTDOC 052416CD funded by VRIDEI-USACH.